# Distributed acoustic sensing for ocean applications

### v2

@@ II.6 Data acquisition
The preceding analysis indicates that setting up a DAS system requires defining a few acquisition parameters, such as the channel distance, the pulse repetition rate, the pulse width or pulse length, and the gauge length.
@@
For unambiguous measurements, a new pulse is emitted only after the previous pulse has traveled along the fiber and returned to the transmission point. Hence, the pulse repetition frequency in a DAS acquisition system depends on the length of the fiber $f_{r}=1/\tau_{\text{max}}=c_{n}/(2L_{f})$. For example, for a 100 km optical fiber cable with a typical refractive index of ca $n\approx 1.46$, $f_{r}$ cannot exceed 1 kHz.
@@
Figure 4 illustrates the timescales of a typical DAS dataset as defined by the sampling frequency $f_{s}$ and the pulse repetition frequency $f_{r}$ [Hartog2017, , Fig. 6.1]. The output of a DAS system is structured in a two-dimensional format $E_{b}(\tau,t)$, where $\tau$ is the fast-time and $t$ is the slow-time. Measurements along the fast-time correspond to the spatial dimension along the fiber, $z=\tau c_{n}/2$, and are commonly downsampled to the channel spacing $L_{h}$. In contrast, measurements along the slow-time correspond to the temporal dimension of the dataset. Hence, the pulse repetition frequency defines the maximum observable acoustic frequency, $f_{a}^{\text{max}}<=f_{r}/2$. Note that the two-dimensional data arrangement resembles the typical data arrangement in synthetic aperture systems for radar [sar2005, , Ch. 4.6.1] or sonar applications.sas2011 In active sensing with synthetic aperture systems, though, time and space are accessed by the probe pulse and the platform motion, respectively, hence fast-time corresponds to the temporal and slow-time to the spatial dimension.
@@
The laser pulse duration $T_{p}$, referred to as pulse width, determines the pulse length $L_{p}=T_{p}c_{n}$. The longer the pulse duration the higher the SNR, but the lower the spatial resolution as the fiber section within a pulse length acts like a single sensor. A pulse length is typically longer than the channel distance, which results in dependent channel measurements within the pulse length. Typical pulse widths are 10–1000 ns, corresponding to a pulse length $L_{p}\approx$ 2–200 m for a fiber with a typical refractive index of ca $n\approx 1.46$.
@@
The gauge length $L_{g}$ determines the overlap between the backscattered signals originating from different locations along the fiber, which are mixed at the interrogator for differential phase measurements. Consequently, the gauge length should be larger than the pulse length for the mixed backscattered signals to be independent of the pulse coherence. Additionally, the gauge length determines the resolution of the axial strain measurement along the fiber; see Eq. (19). The greater the gauge length compared to the pulse length, the more linear the relation between differential phase and strain [Hartog2017, , Fig. 6.29]. On the other hand, the impinging wavefield should remain spatially coherent within a gauge length for meaningful measurements. To fulfill these competing requirements, the gauge length is commonly set equal to the pulse length.
@@
Figure 5: (color online) Fourier pairs of distributed sensing parameters: Space (left) and wavenumber (right). (a) Fiber aperture with a total length of $L_{f}$, (b) spatial sampling at channel spacing $L_{h}$, (c) pulse aperture of length $L_{p}$, (d) gauge aperture of length $L_{g}$, and (e) total distributed sensing function resulting from sampling the total fiber aperture at the channel spacing (multiplication) and averaging over the pulse and gauge length (convolution).
@@
Figure 6: (color online) DAS directivity when setting the gauge length $L_{g}$ equal to the pulse length $L_{p}$ as a function of (a) the wavenumber along the fiber axis $k_{z}$ and (b) the direction of the incident angle $\sin\theta$. (c) The directional sensitivity of the axial strain.
@@ III Distributed vs array processing
Figure 5 illustrates the aperture sampling and the directivity of the distributed sensing modality. The optical fiber with length $L_{f}$ is represented as a rectangular function $\mathit{f}(z)$ in one-dimensional space $z$ and its Fourier transform $\widetilde{F}(k_{z})$ is the sinc function in the wavenumber domain $k_{z}=k_{a}\sin\theta/(2\pi)$; Fig. 5(a1)–(a2),
$\mathit{f}(z)=\Pi(\frac{z}{L_{f}})=\begin{cases}1,\,&\text{for}\quad\lvert z\rvert\leq\frac{L_{f}}{2}\\ 0,\,&\text{otherwise},\end{cases}$ (22)
$\widetilde{F}(k_{z})=\mathcal{F}(\mathit{f}(z))=\mathrm{sinc}(L_{f}k_{z})=\frac{\sin(\pi L_{f}k_{z})}{\pi L_{f}k_{z}}.$
The differential phase of the backscattered light from a laser pulse traveling along the optical fiber is sampled at $n$ channels distributed at regular spacings $L_{h}$ along the fiber, such that $L_{f}=nL_{h}$. This sampling is represented by a Dirac comb function $h(z)$ with period $L_{h}$ and its Fourier transform is another Dirac comb function $\widetilde{H}(k_{z})$ with period $1/L_{h}$; Fig. 5(b1)–(b2),
$h(z)=\mathrm{III}_{L_{h}}(z)=\sum\limits_{i=-\infty}^{\infty}\delta(z-iL_{h}),$ (23)
$\widetilde{H}(k_{z})=\mathcal{F}(h(z))=\mathrm{III}_{1/L_{h}}(k_{z})=\sum\limits_{i=-\infty}^{\infty}\delta(k_{z}-i\frac{1}{L_{h}}),$
@@
As the probe pulse propagates along the fiber, it creates a moving average filter for the backscattered field over the pulse length that spans, potentially, multiple channels, $L_{p}=mL_{h}$ with $m\geq 1$. Using a pulse length at least equal to the spatial sampling distance determined by the channel spacing results in a densely populated sensing array as the ones commonly used for synthetic apertures in sonar and radar applications.Rolt1992 The pulse aperture (3) is represented here for illustration simplicity by a rectangular function $\mathit{w}(z)$ with Fourier transform $\widetilde{W}(k_{z})$; Fig. 5(c1)–(c2),
$\mathit{w}(z)=\Pi(\frac{z}{L_{p}})=\begin{cases}1,\,&\text{for}\quad\lvert z\rvert\leq\frac{L_{p}}{2}\\ 0,\,&\text{otherwise},\end{cases}$ (24)
$\widetilde{W}(k_{z})=\mathcal{F}(\mathit{w}(z))=\mathrm{sinc}(L_{p}k_{z})=\frac{\sin(\pi L_{p}k_{z})}{\pi L_{p}k_{z}}.$
Estimating strain over a finite gauge length introduces another spatial filter to the measurement of external vibrations; see (19). The gauge length can be represented as a rectangular aperture $\mathit{g}(z)$ with Fourier transform $\widetilde{G}(k_{z})$; Fig. 5(d1)–(d2),
$\mathit{g}(z)=\Pi(\frac{z}{L_{g}})=\begin{cases}1,\,&\text{for}\quad\lvert z\rvert\leq\frac{L_{g}}{2}\\ 0,\,&\text{otherwise},\end{cases}$ (25)
$\widetilde{G}(k_{z})=\mathcal{F}(\mathit{g}(z))=\mathrm{sinc}(L_{g}k_{z})=\frac{\sin(\pi L_{g}k_{z})}{\pi L_{g}k_{z}}.$
Since the differential phase measurements are unambiguously related to strain from external vibrations only for $L_{g}\geq L_{p}$ (see Sec. II.4), it is the gauge length that determines the final spatial resolution for DAS measurements.
The resulting distributed aperture $d(z)$ is constructed from sampling the optical fiber at channel spacing, $(f\times h)(z)$, and convolving the result with the pulse. Due to averaging over the gauge length, the output is convolved with (25). Combining (22)–(25) in either spatial or wavenumber domain yields
@@
as depicted in Fig. 5(e1)–(e2)
Setting $L_{g}=L_{p}$ results in a directivity that is determined by
$(\widetilde{W}\times\widetilde{G})(k_{z})\!=\!\mathrm{sinc}(L_{p}k_{z})\mathrm{sinc}(L_{g}k_{z})\!=\!\mathrm{sinc}^{2}(L_{g}k_{z}),$ (27)
as depicted in Fig. 6(a) or, equivalently, in Fig. 6(b) rescaling the wavenumber axis by the wavelength $\lambda_{a}$ to indicate the directional dependence $\sin\theta$. The axial strain estimate has a directional sensitivity with respect to the incident angle $\theta$, which is proportional to the square of the sine of the incident angle (19), as shown in Fig. 6(c). Therefore, the higher the frequency (smaller wavelengths) or the longer the gauge length, the more directive the DAS response to external vibrations becomes. However, the more directive the DAS response is, the less sensitive the measurement is to external fields due to the directional response of the axial strain. Low frequencies such that $\lambda_{a}\geq L_{g}$ are detected with higher sensitivity.
Figure 7: (color online) Deployment geometry of the north cable. The fiber portion corresponding to the data analyzed herein is indicated.
Table 2: Dataset parameters.
Parameter Dataset
description formula unit A, C B
refractive index $n_{f}$ 1.4682 1.4682
@@
number of slow-time samples $n_{t}$ 12000 15000
@@ IV Sound source detection
The data analysis presented in this section is based on material provided by the Ocean Observatories Initiative (OOI), a major facility fully funded by the National Science Foundation under Cooperative Agreement No. 1743430, and the Woods Hole Oceanographic Institution OOI Program Office.dataOOI The analysis is based on datasets collected by the OOI Regional Cabled Array deployed offshore Central Oregon in November 2021. This region was covered in the Cascadia initiative with 60 Ocean bottom seismometers. toomey2014cascadia The data and associated metadata are publicly available by the University of Washington.
The analysis herein is based on three datasets. Datasets A: North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T022302Z and B: North-C2-HF-P1kHz-GL30m-Sp2m-FS500Hz_2021-11-03T015731Z feature acoustic signals that are attributed to whale calls and ship noise, respectively.wilcock2023distributed Marine mammal vocalizations and ship noise are loud acoustic sources that are detectable over long distances. The emanating wavefields insonify a substantial length of the fiber optic cable and can serve as sources of opportunity for inferring environmental parameters, such as sediment thickness and density.kuna2021seismic
Dataset C features lower frequency acoustic signals in the water column due to a distant earthquake. Such signals are known as T waves, i.e., Tertiary waves, as they arrive after P or S waves.okal2008generation The seismic waves from the earthquake initially propagate through the earth causing high frequencies to be filtered out and then couple into the water column on sloping bathymetry near the earthquake.degroot2001excitation The resulting T waves propagating as acoustic waves in the water column tend to be low frequency with a peak energy at frequencies below 10 Hz. To appreciate the spatio-temporal distribution of T waves, dataset C concatenates the data in four successive files, namely North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093202Z, North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093302Z, North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093402Z, and North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093502Z.
@@
Figure 8: (color online) (a) Recorded data in strain units (b) corresponding f-k representation indicating the dominant very low-frequency energy of ocean and seismic waves. (c) The same data after high-pass filtering featuring acoustic wavefronts propagating along the fiber and (d) the corresponding f-k representation.
Figure 9: (a) Wavenumber-time representation of the data in Fig. 8(c) and (b) the average response across time with notches at multiples of $k=1/L_{g}$.
Figure 10: (color online) Estimates of (a) the frequency $f_{a}$ (29) and (b) the angle of incidence $\theta$ (30) of the acoustic wavefield sensed by the optical fiber cable.
Figure 11: (color online) (a) The phase and (b) the spatial phase gradient of the analytical signal in Fig. 8(a) corresponding to the earliest arrival. The corresponding values for a fitted simulation example, assuming either (c), (d) plane wave propagation or (e), (f) spherical wave propagation.
Figure 12: (color online) Geometry of spherical wave propagation for an omnidirectional source at range $r$ from the point of closest approach $z_{\rm pca}$ and corresponding incidence angle as a function of axial distance $z$.
@@ IV.1 Whale calls
Dataset A was collected along the transmit fiber in the north cable on 4/11/2021. The duration of the recording is 60 s sampled at 200 Hz. The channel spacing is $L_{h}=2.0419$ m and the gauge length is set to $L_{g}=25L_{h}\approx L_{p}\approx 50$ m.
Figure 8(a) shows the time series recorded over 10000 channels (channels 10000–20000) after de-trending, i.e., subtracting the mean per recording and transforming the differential phase measurement to strain (20). Applying the two-dimensional Fourier transform to the recorded data, the frequency-wavenumber (f-k) representation in Fig. 8(b) is obtained, which shows that the recorded signal is dominated by very low-frequency ocean and seismic waveswilliams2019distributed . In Fig. 8(b), positive wavenumbers represent wave propagation from the ocean toward the shore and vice versa. Signals at the frequency range 0.2–1 Hz are attributed to seismic Rayleigh/Scholte waves with phase velocity in the range 200–1000 m/s and more energy propagating from shore toward the ocean; see Sec. VI.1. The lower-frequency signals ($<0.1$ Hz) represent ocean surface gravity and infragravity waves, with phase velocity ca. 20 m/s and more energy propagating from the ocean toward the shore; see Sec. VI.2.
To remove the dominant very low-frequency signals, the data are filtered with a digital eight-order Butterworth high-pass filterscipy with a cut-off frequency of 15 Hz. The high-pass filtered data in time-space and f-k representation, depicted in Figs. 8(c) and (d) respectively, indicate narrowband (15–20 Hz) acoustic arrivals that repeat at regular intervals (every ca. 20 s) and are classified as finback whale calls.wilcock2023distributed ; watkins1987 The f-k spectrum of the filtered data in Fig. 8(d) shows that these signals propagate in both directions along the cable at an apparent phase velocity of ca. 1500 m/s, consistent with an acoustic wave in the water column.
Figure 8(d) indicates that there are regular notches along the wavenumber dimension. To show the wavenumber pattern clearly, Fig. 9(a) shows a representation of the data in Fig. 8(c) in the time-wavenumber domain, i.e., applying the Fourier transform only along the spatial dimension, and Fig. 9(b) shows the average across time. The notches along the wavenumber dimension occur at multiples of $k=\pm 1/L_{g}\approx\pm 0.02$ m-1 in accordance to the gauge length beampattern; see Fig. 5(d2).
Applying the Hilbert transformOppenheimBook ; scipy on the real-valued two-dimensional dataset $d(z,t)$ in Fig. 8(c) provides the complex-valued analytic expression for the estimated strain (19) $\widehat{\epsilon}_{zz}(z,t)=d(z,t)*\left[\delta(t)+j/\pi t\right]$ such that the instantaneous phase is recovered, i.e.,
$\phi_{a}(z,t)=\angle{\widehat{\epsilon}_{zz}(z,t)}=\omega_{a}t-k_{a}\sin\theta z.$ (28)
The temporal gradient of the unwrapped phase (28) at each location $z_{c}$ is proportional to the acoustic signal frequency as
@@
Whereas, the spatial gradient of the unwrapped phase (28) at each time instant $t_{c}$ is proportional to the acoustic wavenumber and the direction of arrival $\theta$ defined in Fig. 3 as
$\Delta_{z}\phi_{a}(z,t_{c})=-k_{a}\sin\theta\Delta z=-\frac{2\pi f_{a}}{c_{a}}\sin\theta L_{h}\quad\Rightarrow$

### v3

@@ II.6 Data acquisition
The preceding analysis indicates that setting up a DAS system requires defining a few acquisition parameters, such as channel spacing, pulse repetition rate, pulse width or pulse length, and gauge length.
@@
For unambiguous measurements, a new pulse is emitted only after the previous pulse has traveled along the fiber and returned to the transmission point. Hence, the pulse repetition frequency in a DAS acquisition system depends on the length of the fiber $f_{r}=1/\tau_{\text{max}}=c_{n}/(2L_{f})$. For example, for a 100 km optical fiber cable with a typical refractive index of $n\approx 1.46$, $f_{r}$ cannot exceed 1 kHz.
@@
Figure 5 illustrates the timescales of a typical DAS dataset as defined by the sampling frequency $f_{s}$ and the pulse repetition frequency $f_{r}$ [Hartog2017, , Fig. 6.1]. The output of a DAS system is structured in a two-dimensional format $E_{b}(\tau,t)$, where $\tau$ is the fast-time and $t$ is the slow-time. Measurements along the fast-time correspond to the spatial dimension along the fiber, $z=\tau c_{n}/2$, and are commonly downsampled to the channel spacing $L_{h}$.
Measurements along the slow-time correspond to the temporal dimension of the dataset [Fig. 5]. The pulse repetition frequency defines the maximum observable acoustic frequency, $f_{a}^{\text{max}}<=f_{r}/2$. The two-dimensional data arrangement resembles the typical data arrangement in synthetic aperture systems for radar [sar2005, , Ch. 4.6.1] or sonar applications.sas2011 In active sensing with synthetic aperture systems, though, time and space are accessed by the probe pulse and the platform motion, respectively, hence, fast-time corresponds to the temporal and slow-time to the spatial dimension.
@@
The laser pulse duration $T_{p}$, or pulse width, determines the pulse length $L_{p}=T_{p}c_{n}$. The longer the pulse duration, the higher the SNR, but the lower the spatial resolution, as the fiber section within a pulse length acts like a single sensor. A pulse length is typically longer than the channel distance, resulting in dependent channel measurements within the pulse length. Typical pulse widths are 10–1000 ns, corresponding to a pulse length $L_{p}\approx$ 2–200 m for a fiber with a typical refractive index of $n\approx 1.46$.
@@
The gauge length $L_{g}$ determines the overlap between the backscattered signals from adjacent locations along the fiber, which are mixed at the interrogator for differential phase measurements. Consequently, the gauge length should be larger than the pulse length for the mixed backscattered signals to be independent of the pulse coherence. Additionally, the gauge length determines the resolution of the axial strain measurement along the fiber; see Eq. (25). The greater the gauge length compared to the pulse length, the more linear the relation between differential phase and strain [Hartog2017, , Fig. 6.29]. On the other hand, the impinging wavefield should remain spatially coherent within a gauge length for meaningful measurements. To fulfill these competing requirements, the gauge length is commonly set equal to the pulse length.spica2023pubdas ; dataOOI
@@
Figure 6: (color online) Fourier pairs of distributed sensing parameters: Space (left) and wavenumber (right). (a) Fiber aperture with a total length of $L_{f}$, (b) spatial sampling at channel spacing $L_{h}$, (c) pulse aperture of length $L_{p}$, (d) gauge aperture of length $L_{g}$, and (e) total distributed sensing function resulting from sampling the total fiber aperture at the channel spacing (multiplication) and averaging over the pulse and gauge length (convolution). For unambiguous measurements, $L_{f}\gg L_{g}\geq L_{p}\geq L_{h}$.
@@
Figure 7: (color online) DAS directivity when gauge length is equal to pulse length, $L_{g}$$L_{p}$, versus (a) the wavenumber along the fiber axis $k_{z}$ and (b) the direction of the incident angle $\sin\theta$. (c) The directional sensitivity of the axial strain.
@@ III Distributed vs array processing
Figure 6 illustrates the aperture sampling and the directivity of the distributed sensing modality. The optical fiber with length $L_{f}$ is represented as a rectangular function $\mathit{f}(z)$ in one-dimensional space $z$ and its Fourier transform $\widetilde{F}(k_{z})$ is the sinc function in the wavenumber domain $k_{z}=k_{a}\sin\theta$; Fig. 6(a1)–(a2),
[…]
$\widetilde{F}(k_{z})=\mathcal{F}(\mathit{f}(z))=\mathrm{sinc}(k_{z}L_{f}/2)=\frac{\sin(k_{z}L_{f}/2)}{k_{z}L_{f}/2}.$
The differential phase of the backscattered light from a laser pulse traveling along the optical fiber is sampled at $n$ channels distributed at regular spacings $L_{h}$ along the fiber, such that $L_{f}=nL_{h}$. This sampling is represented by a Dirac comb function $h(z)$ with period $L_{h}$ and its Fourier transform is another Dirac comb function $\widetilde{H}(k_{z})$ with period $2\pi/L_{h}$; Fig. 6(b1)–(b2),
$h(z)=\sum\limits_{i=-\infty}^{\infty}\delta(z-iL_{h}),$ (29)
$\widetilde{H}(k_{z})=\mathcal{F}(h(z))=\sum\limits_{i=-\infty}^{\infty}\delta(k_{z}-i\frac{2\pi}{L_{h}}),$
@@
The fiber is probed by a narrowband light pulse defined by the window in Eq. (3), which is represented here for illustration simplicity by a rectangular function $\mathit{w}(z)$ with Fourier transform $\widetilde{W}(k_{z})$; Fig. 6(c1)–(c2),
[…]
$\widetilde{W}(k_{z})=\mathcal{F}(\mathit{w}(z))=\mathrm{sinc}(k_{z}L_{p}/2)=\frac{\sin(k_{z}L_{p}/2)}{k_{z}L_{p}/2}.$
Estimating strain over a finite gauge length introduces another spatial filter to the measurement of external vibrations; see Eq. (25). The gauge length can be represented as a rectangular aperture $\mathit{g}(z)$ with Fourier transform $\widetilde{G}(k_{z})$; Fig. 6(d1)–(d2),
[…]
$\widetilde{G}(k_{z})=\mathcal{F}(\mathit{g}(z))=\mathrm{sinc}(k_{z}L_{g}/2)=\frac{\sin(k_{z}L_{g}/2)}{k_{z}L_{g}/2}.$
The resulting distributed aperture $d(z)$ is constructed from sampling the optical fiber at channel spacing, $(f\times h)(z)$, and convolving the result with the pulse. Due to averaging over the gauge length, the output is convolved with the gauge length function in Eq. (31). Combining Eqs. (28)–(31) in either spatial or wavenumber domain yields
@@
as depicted in Fig. 6(e1)–(e2).
The pulse length potentially spans multiple channels, $L_{p}=mL_{h}$ with $m\geq 1$. Using a pulse length at least equal to the spatial sampling distance determined by the channel spacing results in a densely populated sensing array, i.e., an array without gaps between the sensors. Such dense sensing attenuates effectively the grating lobes introduced by the spatial sampling; see Fig. 6(e2).
Since differential phase measurements are unambiguously related to strain from external vibrations only for $L_{g}\geq L_{p}$ (see Sec. II.4), it is the gauge length that determines the final spatial resolution for DAS measurements. Setting $L_{g}=L_{p}$ results in a directivity that is determined by,
$(\widetilde{W}\times\widetilde{G})(k_{z})$ $=\mathrm{sinc}(k_{z}L_{p}/2)\mathrm{sinc}(k_{z}L_{g}/2)$ (33)
$=\mathrm{sinc}^{2}(k_{z}L_{g}/2),$
as depicted in Fig. 7(a) or, equivalently, in Fig. 7(b) rescaling the wavenumber axis by $\lambda_{a}/(2\pi)$ to indicate the directional dependence $\sin\theta$. The higher the frequency (smaller wavelengths) or the longer the gauge length, the more directive the DAS response to external vibrations becomes. However, the axial strain estimate has a directional sensitivity with respect to the incident angle $\theta$, which is proportional $\sin^{2}\theta$; see Eq. (25) and Fig. 7(c). Therefore, the more directive the DAS response is, the less sensitive the measurement is to external fields due to the directional response of the axial strain [Fig. 7]. Low frequencies with $\lambda_{a}\geq L_{g}$ are detected with higher sensitivity.
Figure 8: (color online) Deployment geometry of the north cable. The fiber portion corresponding to the data analyzed herein is indicated. Dataset C corresponds to the whole cable.
Table 2: List of acquisition parameters and their respective values for each of the datasets used for the analysis.
Parameter
description symbol unit value
Dataset
A, C B
refractive index $n_{f}$ - 1.4682 1.4682
@@
number of slow-time samples $n_{t}$ - 12000 15000
@@ IV Sound source detection
The data analysis presented in this section is based on material provided by the Ocean Observatories Initiative (OOI). The analysis is based on datasets collected by the OOI Regional Cabled Array deployed offshore Central Oregon in November 2021. This region was covered in the Cascadia initiative, which included 60 Ocean bottom seismometers. toomey2014cascadia The data and associated metadata are publicly available by the University of Washington.dataOOI
The analysis herein is based on three datasets. Datasets A (North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T022302Z) and B (North-C2-HF-P1kHz-GL30m-Sp2m-FS500Hz_2021-11-03T015731Z) feature acoustic signals that are attributed to whale calls and ship noise, respectively.wilcock2023distributed Marine mammal vocalizations and ship noise are loud acoustic sources that are detectable over long distances. The emanating wavefields insonify a substantial length of the fiber-optic cable and can serve as sources of opportunity for inferring environmental parameters, such as sediment thickness and density.kuna2021seismic
Dataset C features lower frequency acoustic signals in the water column due to a distant earthquake. Such signals are known as T waves, i.e., Tertiary waves, as they arrive after primary (pressure) P waves and secondary (shear) S waves.okal2008generation The seismic waves from the earthquake initially propagate through the earth causing high frequencies to be filtered out and then couple into the water column on sloping bathymetry near the earthquake.degroot2001excitation The resulting T waves propagating as acoustic waves in the water column tend to be low frequency with a peak energy below 10 Hz. To capture the full temporal extent of T waves, dataset C concatenates the data in four successive files, namely North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093202Z, North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093302Z, North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093402Z, and North-C1-LR-P1kHz-GL50m-Sp2m-FS200Hz_2021-11-04T093502Z.
@@
Figure 9: (color online) (a) Recorded data in strain units as a function of distance $z$ from the interrogator and slow time $t$ and (b) corresponding f-k representation indicating the dominant very low-frequency energy of ocean and seismic waves. (c) The same data after high-pass filtering featuring acoustic wavefronts propagating along the fiber and (d) the corresponding f-k representation.
Figure 10: (a) Wavenumber-time representation of the data in Fig. 9(c) and (b) the average response across time with notches at multiples of $k=2\pi/L_{g}=2\pi/51=0.123{\rm rad/m}$.
Figure 11: Estimates of (a) the frequency $f_{a}$ in Eq. (35) and (b) the angle of incidence $\theta$ in Eq. (36) of the acoustic wavefield featuring whale calls sensed by the optical fiber cable.
Figure 12: (color online) (a) The phase and (b) the spatial phase gradient of the analytical signal in Fig. 9(a) corresponding to the whale call signal at around 10 s in the recording. The corresponding values for a fitted simulation example, assuming either plane wave propagation in (c) and (d) or spherical wave propagation in (e) and (f).
Figure 13: (color online) (a) Geometry of spherical wave propagation for an omnidirectional source at range $r$ from the closest point of approach $z_{\rm cpa}$ and (b) corresponding incident angle as a function of axial distance $z$.
@@ IV.1 Whale calls
Dataset A was collected along the transmit fiber in the north cable on November 4, 2021. The duration of the recording is 60 s sampled at 200 Hz. The channel spacing is $L_{h}=2.0419$ m and the gauge length is set to $L_{g}=25L_{h}\approx L_{p}\approx 50$ m.
Figure 9(a) shows the DAS data recorded over 10000 channels (channels 10000–20000) after de-trending, i.e., subtracting the mean per recording and transforming the differential phase measurement to strain, Eq. (26). Applying the two-dimensional Fourier transform to the recorded data, the frequency-wavenumber (f-k) representation in Fig. 9(b) is obtained, which shows that the recorded signal is dominated by very low-frequency ocean and seismic waves.williams2019distributed In Fig. 9(b), positive wavenumbers represent wave propagation from the ocean toward the shore and vice versa. Signals at the frequency range 0.2–1 Hz are attributed to seismic Rayleigh/Scholte waves with phase velocity in the range 200–1000 m/s and more energy propagating from shore toward the ocean; see Sec. VI.1. The lower-frequency signals (<0.1<0.1 Hz) represent ocean surface gravity and infragravity waves, with phase velocity ca. 20 m/s and more energy propagating from the ocean toward the shore; see Sec. VI.2.
To remove the dominant very low-frequency signals, the data are filtered with a digital eight-order Butterworth high-pass filterscipy with a cut-off frequency of 15 Hz. The high-pass filtered data in time-space and f-k representation, depicted in Figs. 9(c) and (d) respectively, indicate narrowband (15–20 Hz) acoustic arrivals that repeat at regular intervals (every ca. 20 s) and are classified as fin whale (Balaenoptera physalus) calls.wilcock2023distributed ; Goestchel2025fin ; watkins1987 The f-k spectrum of the filtered data in Fig. 9(d) shows that these signals propagate in both directions along the cable at an apparent phase velocity $c_{a}=2\pi f/k\approx 1500$ m/s, consistent with an acoustic wave in the water column.
Figure 9(d) indicates that there are regular notches along the wavenumber dimension. To show the wavenumber pattern clearly, Fig. 10(a) shows a representation of the data in Fig. 9(c) in the time-wavenumber domain, by applying the Fourier transform only along the spatial dimension, and Fig. 10(b) shows the average across time. The notches along the wavenumber dimension occur at multiples of $k=\pm 2\pi/L_{g}\approx\pm 0.125$ rad/m in accordance to the gauge length beampattern; see Fig. 6(d2).
Applying the Hilbert transformOppenheimBook ; scipy on the real-valued two-dimensional dataset $d(z,t)$ in Fig. 9(c) provides the complex-valued analytic expression for the strain in Eq. (25), $\widetilde{\epsilon}_{zz}(z,t)=d(z,t)*\left[\delta(t)+j/\pi t\right]$ such that the instantaneous phase is recovered, i.e.,
$\phi_{a}(z,t)=\angle{\widehat{\epsilon}_{zz}(z,t)}=\omega_{a}t-k_{a}z\sin\theta.$ (34)
The temporal gradient of the unwrapped phase in Eq. (34) at each location $z_{c}$ is proportional to the acoustic signal frequency as
@@
Whereas, the spatial gradient of the unwrapped phase in Eq. (34) at each time instant $t_{c}$ is proportional to the acoustic wavenumber and the incident angle $\theta$ defined in Fig. 4 as
$\Delta_{z}\phi_{a}(z,t_{c})=-k_{a}\Delta z\sin\theta=-\frac{2\pi f_{a}}{c_{a}}L_{h}\sin\theta\quad\Rightarrow$